# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

### v1

@@ 4.2 Comparisons to State-of-the-Art Detectors
Table 4: Ablation study of components in FlexEvent. FlexFuser denotes the adaptive event-frame fusion module (cf. Sec.3.2). FAL denotes the frequency-adaptive learning mechanism (cf. Sec. 3.3). The reported are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL.
Modality FAL FlexFuser Frequency (Hz) Average
$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
E $53.2\%$ $54.0\%$ $53.5\%$ $52.0\%$ $49.4\%$ $45.9\%$ $38.8\%$ $22.9\%$ $46.2\%$
$54.6\%$ $54.9\%$ $54.9\%$ $54.3\%$ $53.3\%$ $50.7\%$ $44.6\%$ $30.4\%$ $49.7\%$
E + F $\blacklozenge$ $54.9\%$ $57.3\%$ $57.7\%$ $57.8\%$ $57.2\%$ $56.1\%$ $53.7\%$ $48.3\%$ $55.4\%$
$\mathbf{58.0\%}$ $59.6\%$ $\mathbf{60.0\%}$ $59.6\%$ $59.0\%$ $57.6\%$ $54.8\%$ $49.2\%$ $57.2\%$
$57.4\%$ $\mathbf{60.0\%}$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $\mathbf{50.9\%}$ $\mathbf{57.9\%}$
@@
Comparison to Multimodal Models. We compare FlexEvent with multimodal event-camera object detection methods such as DAGr (Gehrig & Scaramuzza, 2024), CAFR (Cao et al., 2024), and RENet (Zhou et al., 2023), which fuse event data with other sensor inputs, such as RGB frames or depth to improve detection accuracy. While these methods enhance performance over event-only approaches, they struggle to adapt to varying operational frequencies and often exhibit inadequate feature fusion in dynamic environments. Our approach addresses these limitations by dynamically balancing the contributions of event and frame data. As a result, we achieve superior performance, such as a $47.4\%$ mAP on DSEC-Detection as shown in Tab. 2, outperforming CAFR ($38.0\%$) and RENet ($29.4\%$). This flexible combination of event and frame data, along with its ability to generalize across different temporal resolutions, enables our method to excel in high-frequency detection scenarios, surpassing state-of-the-art methods.
Comparisons Across Different Categories. We evaluate the performance of FlexEvent across various object categories, including cars, pedestrians, and large vehicles. As shown in Tab. 2, we consistently outperform other methods, achieving $59.3\%$ mAP for cars, compared to $49.9\%$ for CAFR (Cao et al., 2024). This highlights its effectiveness in detecting larger, fast-moving objects, while also surpassing CAFR in pedestrian categories. Existing Event-based detectors struggle with detecting smaller and stationary objects due to a lack of pixel intensity changes, and fusion methods over-rely on frame data, which lacks temporal resolution. Our approach addresses these issues with adaptive fusion and temporal refinement, ensuring accurate detection across different object types and motion dynamics. This versatility reinforces its superiority over state-of-the-art methods.
@@
Generalization on High-Frequency Data. A key contribution of FlexEvent is its ability to generalize across various operational frequencies, particularly in high-frequency scenarios. We evaluate this by testing detection performance at different temporal offsets, $\frac{i}{n}\Delta T$, where $n=10$, $i=0,...,10$, and $\Delta T=50$ ms. Ground truth labels are generated by linearly interpolating object positions between frames for consistent evaluation. In this setting, event-based methods are tested across multiple time durations, while event-frame fusion methods process one RGB frame followed by event data of varying time durations. The comparison results are shown in  Fig. 4. Most existing methods, such as RVT and SAST, struggle at higher frequencies due to their fixed temporal intervals and limited ability to capture fast scene changes. In contrast, our approach achieves $56.5\%$ mAP at $90$ Hz and $50.9\%$ at $180$ Hz. This improvement demonstrates that our method excels in dynamic, rapidly changing environments where accurate detection is critical for safety and reliability.
Qualitative Assessment. We provide qualitative comparisons between FlexEvent and other state-of-the-art methods under different event operation frequencies, as shown in Fig. 5 and Fig. 6. Unlike RVT and DAGr, which miss critical object details, our model consistently detects objects with high accuracy, even in challenging cases involving fast-moving vehicles and occluded pedestrians. For instance, in Scene 2 of Fig. 5, RVT fails to detect a pedestrian due to insufficient event data, while our approach successfully identifies the pedestrian by leveraging both frames and high-frequency event data. Similarly, in Scene 4, DAGr struggles with the rapid motion of a large vehicle, leading to inaccurate predictions, whereas our approach ensures precise object localization. In the upper instance of Fig. 6, RVT fails to detect a pedestrian at high frequency due to insufficient event data. In contrast, FlexEvent successfully identifies the pedestrian by leveraging the frame-based information. These qualitative findings confirm that our model excels not only in quantitative metrics but also in real-world performance.
Figure 6: Qualitative comparisons of FlexEvent and RVT (Gehrig & Scaramuzza, 2023) under different event operation frequencies. Our approach demonstrates a strong robustness under both low- and high-frequency scenarios. Best viewed in colors. See the Appendix for more examples.
@@
Table 5: Ablation study of hyperparameter configurations in the FlexEvent frameworks. $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
 $\mathbf{\tau}^{\text{car}}$  $\mathbf{\tau}^{\text{ped}}$ $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) Average
$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
$0.6$ $0.3$ $10$ $0.8$ $56.5\%$ $55.9\%$ $56.7\%$ $57.2\%$ $57.1\%$ $56.7\%$ $54.5\%$ $49.2\%$ $55.5\%$
$0.6$ $0.3$ $10$ $0.6$ $56.7\%$ $57.2\%$ $57.7\%$ $57.9\%$ $57.7\%$ $57.0\%$ $54.3\%$ $47.0\%$ $55.7\%$
$0.6$ $0.3$ $8$ $0.6$ $56.3\%$ $58.5\%$ $58.8\%$ $59.1\%$ $58.8\%$ $58.4\%$ $56.2\%$ $\mathbf{51.2\%}$ $57.2\%$
$0.6$ $0.3$ $6$ $0.6$ $57.3\%$ $59.4\%$ $59.7\%$ $59.9\%$ $59.3\%$ $58.5\%$ $55.7\%$ $48.8\%$ $57.3\%$
$0.6$ $0.6$ $6$ $0.6$ $\mathbf{57.4\%}$ $\mathbf{60.0}\%$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $50.9\%$ $\mathbf{57.9\%}$
$0.8$ $0.8$ $6$ $0.6$ $56.6\%$ $58.7\%$ $59.1\%$ $58.9\%$ $58.4\%$ $57.4\%$ $55.6\%$ $50.2\%$ $56.9\%$
4.3 Ablation studies
Component Analysis. We conduct ablation studies by selectively removing key modules: the FlexFuser and FAL mechanisms. As shown in Tab. 4, removing the FAL mechanism results in a significant performance drop, particularly in high-frequency scenarios, underscoring its role in adapting to varying frequencies and generating frequency-adjusted labels. Similarly, omitting the FlexFuser module leads to a marked decrease in mAP, highlighting the importance of adaptive event-frame fusion for accurate detection across different operational frequencies. We also test training with interpolated labels for high-frequency testing, but this approach reduces recall by missing objects that suddenly appear or disappear, making it less effective than FAL.
Hyperparameter Tuning. We tune the hyperparameters of the FAL mechanism, focusing on key settings like the confidence threshold ($\tau$), IoU threshold, and track length for temporal refinement. As shown in Tab. 5, lowering the confidence threshold improves recall but reduces precision, as the model becomes more lenient in detecting objects. Applying overly strict conditions, such as a higher confidence threshold or IoU threshold, lowers recall by filtering out valid detections. The optimal configuration is achieved with $\tau=0.6$ and a track length of 6, balancing precision and recall for both low- and high-frequency conditions. These moderate settings ensure that FlexEvent maintains robust performance and stable detection accuracy across diverse environments.
@@ 5 Conclusion
This paper introduces FlexEvent, an event-camera object detection framework designed to operate across arbitrary frequencies. By combining FlexFuser for adaptive event-frame fusion and FAL for frequency-adaptive learning, we leverage the rich temporal information of event data and the semantic detail of RGB frames to overcome the limitations of existing methods, providing a flexible solution for dynamic environments. Extensive experiments on large-scale datasets show that our approach significantly outperforms state-of-the-art methods, particularly in high-frequency scenarios, demonstrating its robustness and adaptability for real-world applications such as autonomous driving and robotics.
@@
In this appendix, we supplement the following materials to support the findings and conclusions drawn in the main body of this paper.
[…]
Appendix A A Additional Implementation Details
@@ A.1 Datasets
In this work, we develop and validate our proposed approach on the large-scale DSEC dataset (Gehrig et al., 2021b). DSEC serves as a high-resolution, large-scale multimodal dataset designed to capture real-world driving scenarios under various conditions. It combines data from stereo Prophesee Gen3 event cameras with a resolution of $640\times 480$ pixels and FLIR Blackfly S RGB cameras operating at $20$ FPS, enabling high-fidelity capture of dynamic scenes. To align the RGB frames with the event camera data, an infinite-depth alignment process is employed, which involves undistorting, rotating, and re-distorting the RGB images. This alignment ensures that the event data and RGB frames are temporally and spatially synchronized.
Table 6: Summary of key statistics from the event camera object detection datasets used in this work.
Dataset Reference Classes Frames Sequences Class Names
DSEC-MOD (Zhou et al., 2023) $1$ $13,314$ $16$ Car
DSEC-Detection (Tomy et al., 2022) $3$ $52,727$ $41$ Car
@@
In our experiments, we utilize three comprehensive versions of DSEC tailored for object detection: DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023). A summary of the key statistics of these datasets is listed in Tab. 6.
  • DSEC-Det (Gehrig & Scaramuzza, 2024): This version is developed by the original DSEC team and includes annotations generated using the QDTrack multi-object tracker (Fischer et al., 2023; Pang et al., 2021). The annotation process combines automated multi-object tracking with manual refinement to ensure high-quality and accurate detection labels. Compared to the original DSEC dataset, DSEC-Det introduces additional sequences specifically designed to capture complex and dynamic urban environments, featuring crowded pedestrian areas, moving vehicles, and diverse lighting conditions. These challenging scenarios provide a rich testing ground for evaluating object detection algorithms in real-world driving settings. DSEC-Det comprises $60$ sequences and $78,344$ frames, making it the most extensive dataset used in this study. It captures diverse urban scenes with dynamic elements, such as crowded pedestrian areas and moving vehicles. Covering eight object categories relevant to autonomous driving – Car, Pedestrian, Bus, Bicycle, Truck, Motorcycle, Rider, and Train – the dataset provides a robust foundation for training and evaluating object detection models across various driving scenarios. In our experiments on DSEC-Det, to maintain consistency with the experimental setup of previous work DAGr (Gehrig & Scaramuzza, 2024), we report results on two categories: Car and Pedestrian.
  • DSEC-Detection (Tomy et al., 2022): This dataset comprises $41$ sequences with a total of $52,727$ frames. Focusing on three fundamental object categories – Car, Pedestrian, and Large Vehicle – this version emphasizes high-precision annotations for these critical classes in autonomous driving. The initial annotations are generated using the YOLOv5 model (Jocher, 2020) on RGB frames, leveraging its robust performance in real-time object detection. These annotations are then transferred to the corresponding event frames through homographic transformation, ensuring spatial alignment between the two modalities. A subsequent manual refinement process corrects any discrepancies and enhances annotation quality, resulting in a dataset that provides accurate and reliable labels for event-based object detection.
  • DSEC-MOD (Zhou et al., 2023): This dataset extends the object detection capabilities to moving-object detection across diverse urban environments. It includes $16$ sequences containing $13,314$ frames and is specifically focused on the Car category, making it highly suitable for complex detection tasks in varied urban settings, such as intersections, highways, and residential areas. Featuring high-frequency and dense annotations, the dataset provides a valuable resource for evaluating the performance of event-based object detectors under challenging real-world conditions.
These three versions of the DSEC dataset together provide a comprehensive platform for benchmarking and evaluating event-based object detection methods, capturing a wide spectrum of scenarios, object categories, and environmental conditions. Among them, DSEC-Det is the largest, most recent and most comprehensive, annotated and released by the original DSEC authors. Thus, we prioritize it as the primary benchmark for reporting results, ensuring relevance and reliability. DSEC-Detection and DSEC-MOD are datasets used by two recent event-frame fusion methods, CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023). To validate the effectiveness of our method, we also report results on these two datasets.

### v3

@@ 4.1 Experimental Settings
Table 5: Ablation on FlexFuse and FlexTune learning mechanism on DSEC-Det [16]. Symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FlexTune.
Tune Fuse Frequency (Hz) Avg
$\mathbf{20.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
      $53.2$ $52.0$ $49.4$ $45.9$ $38.8$ $22.9$ $43.7$
$54.6$ $54.3$ $53.3$ $50.7$ $44.6$ $30.4$ $48.0$
$\blacklozenge$ $54.9$ $57.8$ $57.2$ $56.1$ $53.7$ $48.3$ $54.7$
$\mathbf{58.0}$ $59.6$ $59.0$ $57.6$ $54.8$ $49.2$ $56.4$
$57.4$ $\mathbf{60.1}$ $\mathbf{59.5}$ $\mathbf{58.8}$ $\mathbf{56.5}$ $\mathbf{50.9}$ $\mathbf{57.2}$
Baselines. To evaluate our method, we compare it against both state-of-the-art event-only and event-frame fusion methods. For event-only detectors, we compare RVT [18], SAST [39], LEOD [55], and SSM [70], retraining them on DSEC-Det [16] following their respective protocols . For event-frame fusion methods, we compare DAGr [16] and HDI-Former [31] on DSEC-Det [16] using results from the original paper and retrain our method on DSEC-Detection [50] and DSEC-MOD [66] using standard settings to compare with CAFR [3] and RENet [66]. For other methods on DSEC-Detection and DSEC-MOD, we reference results reported in the CAFR and RENet papers. For more details, please refer to Appendix Sec. A.2.
Evaluation Metrics. We evaluate object detectors using the mean Average Precision (mAP) as the primary metric, along with AP50, AP75, APS, APM, and APL from the COCO evaluation protocol [35]. These metrics provide a comprehensive assessment of detection performance across different IoU thresholds and object sizes. Please refer to the Appendix Sec. A.3 for more details.
Implementation Details. Our training follows YOLO-X [63], using the standard detection loss consisting of IoU loss, classification loss, and regression loss, plus a lightweight regularizer that balances the contributions of each modality. The model is trained for 100,000100{,}000 iterations with a batch size of $8$ and a sequence length of $11$, using a learning rate of 11e-44. Experiments are conducted on two NVIDIA RTX A50005000 GPUs with 2424GB memory, with the entire training process completed in approximately one day. Due to space limits, more details are in Appendix Sec. A.4.
4.2 Comparisons to State of the Arts
Comparisons with Event-Only Detectors. We compare FlexEvent with state-of-the-art event-only detectors, including RVT [18], SSM [70], SAST [39], and LEOD [55], as shown in Tab. 1. Our model substantially outperforms all these methods, with the gap widening at higher event frequencies. Event-only methods struggle to maintain detection accuracy at higher frequencies because they lack rich semantic cues. In contrast, we overcome these limitations through the FlexFuse module, which integrates RGB data to compensate for the lack of semantic richness in the event stream. By fusing both event and frame data, we excel in complex and dynamic environments, achieving superior detection accuracy where event-only methods fall short.
@@ 4.2 Comparisons to State of the Arts
Comparisons with Multi-Modal Detectors. We compare FlexEvent with multimodal event-camera object detection methods such as DAGr [16], HDI-Former [31], CAFR [3], and RENet [66], which fuse event data with other sensor inputs to improve detection accuracy. The results are shown in Tab. 1, Tab. 2 and Tab. 6. While these methods enhance performance over event-only approaches, they struggle to adapt to varying operational frequencies and often exhibit inadequate feature fusion in dynamic environments. Our approach addresses these limitations by dynamically balancing the contributions of event and frame data. This flexible combination, along with the ability to generalize across different temporal resolutions, enables our method to excel in high-frequency event-based detection scenarios, surpassing state-of-the-art approaches.
@@
Generalization on High-Frequency Data. A key contribution of FlexEvent is its generalization capability across various operational frequencies, particularly in high-frequency scenarios. We evaluate this by testing detection performance at different temporal offsets, $\frac{i}{n}\Delta T$, where $n=10$, i=0,,10i=0,...,10, and $\Delta T=50$ms. Ground truth labels are generated by linearly interpolating object positions between frames following DAGr [16]. In this setting, event-based methods are tested across multiple time durations, while event-frame fusion methods process one RGB frame followed by event data of varying time durations. As shown in Fig. 6, existing methods degrade significantly at higher frequencies due to fixed temporal intervals and limited ability to capture fast scene changes, while our method achieves consistent, strong performance.
Comparisons on Efficiency. In Tab. 5, we compare inference times and parameter counts using an NVIDIA A50005000 2424GB GPU and an AMD EPYC $32$-Core Processor. Although FlexEvent comprises slightly more parameters overall, its speed matches SAST [39] and clearly surpasses DAGr [16] at all tested frequencies. As FlexTune is performed off-line, the high-frequency generalizability is already propagated in the model and introduces no runtime overhead. Our lightweight fusion head contributes only $\sim\!1.5$M parameters, while the event–frame backbone adds a further $\sim\!44$M that is already highly optimized and incurs virtually no extra latency. This confirms our efficiency and real-time suitability.
Qualitative Assessments. We provide visual comparisons of FlexEvent and other state-of-the-art methods under different event operation frequencies, as shown in Fig. 5. Unlike RVT [18] and DAGr [16], which often miss or misinterpret critical objects, our model consistently detects objects with high accuracy, even in challenging cases involving fast-moving vehicles and occluded pedestrians. This robustness is also evident in examples of Fig. 7, where our method detects a pedestrian missed by RVT [18] due to sparse event data at high frequencies. For more detailed analyses, visualizations, and failure cases, please refer to Appendix Sec. C.
Figure 7: Qualitative assessment of FlexEvent and RVT [18] under different operation frequencies.
Figure 8: Ablation on event-frame fusion strategies under different frequencies on DSEC-Det [16].
@@
Table 6: Ablation study on hyperparameters. $\mathbf{\tau}^{\mathrm{c}}$, $\mathbf{\tau}^{\mathrm{p}}$ denotes the confidence threshold for Car and Pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det [16].
 $\mathbf{\tau}^{\mathrm{c}}$  $\mathbf{\tau}^{\mathrm{p}}$ $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) Avg
$\mathbf{20.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
$0.6$ $0.3$ $10$ $0.8$ $56.5$ $57.2$ $57.1$ $56.7$ $54.5$ $49.2$ $55.2$
$0.6$ $0.3$ $10$ $0.6$ $56.7$ $57.9$ $57.7$ $57.0$ $54.3$ $47.0$ $55.1$
$0.6$ $0.3$ 88 $0.6$ $56.3$ $59.1$ $58.8$ $58.4$ $56.2$ $\mathbf{51.2}$ $56.7$
$0.6$ $0.3$ 66 $0.6$ $57.3$ $59.9$ $59.3$ $58.5$ $55.7$ $48.8$ $56.6$
$0.6$ $0.6$ 66 $0.6$ $\mathbf{57.4}$ $\mathbf{60.1}$ $\mathbf{59.5}$ $\mathbf{58.8}$ $\mathbf{56.5}$ $50.9$ $\mathbf{57.2}$
$0.8$ $0.8$ 66 $0.6$ $56.6$ $58.9$ $58.4$ $57.4$ $55.6$ $50.2$ $56.2$
4.3 Ablation Study
Analysis of FlexFuse. We assess the effect of integrating frame data via FlexFuse. As shown in Tab. 5, adding frame information boosts average mAP from $43.7\%$ to $56.4\%$, with greater gains at higher frequencies where event data alone becomes sparse. Our adaptive gating fuses event and frame features more effectively than simple Add, Concat, or vanilla Attention (Fig. 8), achieving the best accuracy across all frequencies and providing robust representations under diverse conditions.
Analysis of FlexTune. In parallel, we assess FlexTune, focusing on scenarios where event data are sparse and fast-evolving. Without it, performance at high frequencies collapses, as seen in the first row of Tab. 5 where mAP at $180$ Hz is only $22.9\%$. However, activating FlexTune raises this score to $30.4\%$, demonstrating the value of iterative self-training and refinement for frequency adaptation. We also test training with interpolated high-frequency labels, but this approach struggles with objects that appear or disappear rapidly, reducing recall of the label. By contrast, FlexTune updates the model with refined, temporally consistent labels, improving detection quality.
Hyperparameter Tuning. We investigate key hyperparameters in the FlexTune mechanism, specifically, the confidence thresholds for cars $\mathbf{\tau}^{\mathrm{c}}$ and pedestrians $\mathbf{\tau}^{\mathrm{p}}$, the IoU threshold for bounding-box association $\mathbf{\tau}^{\text{iou}}$, and the minimum track length for temporal refinement $\mathbf{L}^{\text{track}}$. As shown in Tab. 8, lowering confidence thresholds can improve recall by admitting more detections, but it risks increasing false positives and thus reducing overall precision. Conversely, setting overly strict thresholds, such as a higher IoU requirement or confidence level, might filter out potential detections, lowering recall. An intermediate setting of $\tau=0.6$ for both classes, with a track length of $6$ and an IoU threshold of $0.6$, emerges as the best trade-off, balancing recall and precision across both low- and high-frequency conditions. These moderate, carefully tuned parameters ensure that FlexEvent remains robust and accurate in a range of real-world scenarios.
@@ 5 Conclusion
This paper introduces FlexEvent, an event-camera detection framework designed to operate across varying frequencies. By combining FlexFuse for adaptive event-frame fusion and FlexTune for frequency-adaptive learning, we leverage the rich temporal information of event data and the semantic detail of RGB frames to overcome the limitations of existing methods, providing a flexible solution for dynamic environments. Extensive experiments on large-scale datasets show that our approach significantly outperforms state-of-the-art methods, particularly in high-frequency scenarios, demonstrating its robustness and adaptability for real-world applications.
@@
Appendix A Additional Implementation Details
@@ A.1 Datasets
In this work, we develop and validate our proposed approach on the large-scale DSEC dataset [17]. DSEC serves as a high-resolution, large-scale multimodal dataset designed to capture real-world driving scenarios under various conditions. It combines data from stereo Prophesee Gen3 event cameras with a resolution of $640\times 480$ pixels and FLIR Blackfly S RGB cameras operating at $20$ FPS, enabling high-fidelity capture of dynamic scenes. To align the RGB frames with the event camera data, an infinite-depth alignment process is employed, which involves undistorting, rotating, and re-distorting the RGB images. This alignment ensures that the event data and RGB frames are temporally and spatially synchronized.
In our experiments, we utilize three comprehensive versions of DSEC tailored for object detection: DSEC-Det [16], DSEC-Detection [50], and DSEC-MOD [66]. A summary of the key statistics of these datasets is listed in Tab. 7.
  • DSEC-Det [16]: This version is developed by the original DSEC team and includes annotations generated using the QDTrack multi-object tracker [11, 38]. The annotation process combines automated multi-object tracking with manual refinement to ensure high-quality and accurate detection labels. Compared to the original DSEC dataset, DSEC-Det [16] introduces additional sequences specifically designed to capture complex and dynamic urban environments, featuring crowded pedestrian areas, moving vehicles, and diverse lighting conditions [29]. These challenging scenarios provide a rich testing ground for evaluating object detection algorithms in real-world driving settings. DSEC-Det [16] comprises $60$ sequences and 78,34478{,}344 frames, making it the most extensive dataset used in this study. It captures diverse urban scenes with dynamic elements, such as crowded pedestrian areas and moving vehicles. Covering eight object categories relevant to autonomous driving, Car, Pedestrian, Bus, Bicycle, Truck, Motorcycle, Rider, and Train, the dataset provides a robust foundation for training and evaluating object detection models across various driving scenarios. In our experiments on DSEC-Det [16], to maintain consistency with the experimental setup of previous work DAGr [16], we report results on two categories: Car and Pedestrian.
  • DSEC-Detection [50]: This dataset comprises $41$ sequences with a total of 52,72752{,}727 frames. Focusing on three fundamental object categories – Car, Pedestrian, and Large Vehicle – this version emphasizes high-precision annotations for these critical classes in autonomous driving. The initial annotations are generated using the YOLOv5 model [26] on RGB frames, leveraging its robust performance in real-time object detection. These annotations are then transferred to the corresponding event frames through homographic transformation, ensuring spatial alignment between the two modalities. A subsequent manual refinement process corrects any discrepancies and enhances annotation quality, resulting in a dataset that provides accurate and reliable labels for event-based object detection.
  • DSEC-MOD [66]: This dataset extends the object detection capabilities to moving-object detection across diverse urban environments. It includes $16$ sequences containing 13,31413{,}314 frames and is specifically focused on the Car category, making it highly suitable for complex detection tasks in varied urban settings, such as intersections, highways, and residential areas. Featuring high-frequency and dense annotations, the dataset provides a valuable resource for evaluating the performance of event-based object detectors under challenging real-world conditions.
These three versions of the DSEC dataset together provide a comprehensive platform for benchmarking and evaluating event-based object detection methods, capturing a wide spectrum of scenarios, object categories, and environmental conditions. Among them, DSEC-Det [16] is the largest, most recent, and most comprehensive, annotated, and released by the original DSEC authors. Thus, we prioritize it as the primary benchmark for reporting results, ensuring relevance and reliability. DSEC-Detection [50] and DSEC-MOD [66] are datasets used by two recent event-frame fusion methods, CAFR [3] and RENet [66]. To validate the effectiveness of our method, we also report results on these two datasets.
Table 7: Summary of key statistics from the three event datasets [16, 50, 66] used in this work.
Dataset Classes Frames Sequences Class Names
DSEC-MOD [66] 11 $13,314$ $16$ Car
DSEC-Detection [50] 33 $52,727$ $41$ Car